OpenStudy (amtran_bus):

write the equation of the sine function Amp:4 Period: 5 PI Phase Shift: -PI/5

OpenStudy (anonymous):

y = 4sin(2theta/5 - pi/50 FINDING B: p = 2pi/b 5pi = 2pi/b 5pi/2pi = 1/b 2/5 = b

OpenStudy (amtran_bus):

Thanks, was there a trick to it...I mean finding a number to use to get your -c/b?

OpenStudy (anonymous):

Not really, just plugging in variables to known equations for a sign function period = 2pi/b y = Ampsin(Btheta +/- C) +/- D was that phases shift in the y or x direction?

OpenStudy (amtran_bus):

y

OpenStudy (anonymous):

ooh then its gonna be: \[Y = 4\sin (2\theta \div 5) - \pi \div 5\]

OpenStudy (amtran_bus):

Thanks so much!